### v1

@@ Abstract
Building energy modeling plays a vital role in optimizing the operation of building energy systems by providing accurate predictions of the building’s real-world conditions. In this context, various techniques have been explored, ranging from traditional physics-based models to data-driven models. Recently, researchers are combining physics-based and data-driven models into hybrid approaches. This includes using the physics-based model output as additional data-driven input, learning the residual between physics-based model and real data, learning a surrogate of the physics-based model, or fine-tuning a surrogate model with real data. However, a comprehensive comparison of the inherent advantages of these hybrid approaches is still missing. The primary objective of this work is to evaluate four predominant hybrid approaches in building energy modeling through a real-world case study, with focus on indoor temperature dynamics. To achieve this, we devise three scenarios reflecting common levels of building documentation and sensor availability, assess their performance, and analyse their explainability using hierarchical Shapley values. The real-world study reveals three notable findings. First, greater building documentation and sensor availability lead to higher prediction accuracy for hybrid approaches. Second, the performance of hybrid approaches depend on the type of building room, but the residual approach using a Feedforward Neural Network as data-driven sub-model performs best on average across all rooms. This hybrid approach also demonstrates a superior ability to leverage the physics-based simulation from the physics-based sub-model. Third, hierarchical Shapley values prove to be an effective tool for explaining and improving hybrid models while accounting for input correlations.
@@ 1 Introduction
In the last decade, the operation of buildings accounted for approximately 30% of global energy consumption and 26% of CO2 emissions [18]. Over half of this consumption stems from the operation of building energy systems such as Heating, Ventilation and Air Conditioning (HVAC), as well as the electrical systems. This situation underscores the urgent need for measures to reduce energy consumption in the operation these systems. Fortunately, advancements in digitalization and sensor deployment provide a foundation for using analytical and machine learning tools to optimize the performance of building energy systems. On this basis, Building Energy Models (BEM) can be developed to simulate a building’s real-world condition and predict future behaviour thereby enabling the recommendation of optimal control actions. However, a major challenge remains: the availability of data, particularly in terms of comprehensive building documentation, sensor coverage, and the length of recorded data. This lack of data complicates the reliable and accurate construction of BEMs but also hinders the widespread adoption of energy optimization strategies in buildings.
The Building Energy Model is a digital representation of a building designed for simulating energy and temperature dynamics. It is constructed based on various characteristics of the building, including building geometry, material properties, installed energy systems, and operational inputs such as weather conditions, HVAC operation, and occupancy schedules [11]. To accurately reflect real-world conditions, the BEM is initially calibrated with sensor data to ensure faithful representation of the building and its energy systems. Once calibrated, the BEM can predict building operational performance and indoor temperature evolution across different time scales such as quarter-hourly or hourly resolution. Its comprehensive predictive capabilities make the BEM a versatile tool for various applications in building planning and operations. In retrofitting studies, it serves as an evaluation tool to assess potential measures for improving energy efficiency [36]. In building operations, it provides detailed temperature predictions as basis for control algorithms, enabling the optimization of energy system operations [20].
@@
More recent approaches of building energy modeling focus on data-driven methods, which utilize statistical and machine learning techniques such as Autoregressive Integrated Moving Average (ARIMA), Feedforward Neural Network (FFNN), Long-Short Term Memory (LSTM) and Convolutional Neural Network (CNN) [11]. These methods rely on sensor measurements to establish relationships between defined input and output variables. Commonly modeled relationships include those between building operation data – such as energy consumption or temperature – and the corresponding sensor readings. Because these models are directly applied to data, the underlying mechanisms are not easily accessible or interpretable, leading them to be classified as data-driven models. Unlike physics-based models, data-driven approaches do not require detailed physical building information or the calibration of physical parameters, making them easier to set up as digital twins. However, their performance is highly dependent on the quantity and quality of the available measurement data. Despite this dependency, data-driven models have demonstrated high accuracy in numerous studies and are particularly well-suited for modelling buildings at an urban scale, thanks to their reduced configuration time [29].
@@
Several research studies aim to combine reduced-order physics-based models with data-driven models through two general strategies. The first strategy involves formulating a loss function based on the reduced-order model to train the data-driven model. A prominent example is the use of Physics-informed Neural Network (PINN), where a physics-informed loss function is devised based on RC model temperature dynamics in order to train a FFNN [17, 26, 8]. The second strategy consists of incorporating reduced-order model elements into a data-driven model architecture. A number of works model building dynamics with a State-Space model framework and parameterize state and/or observation equation with FFNN [14], Neural Ordinary Differential Equations [38] or Graph Neural Networks [47]. However, ensuring physical consistency still remains a challenge in aforementioned strategies [12].
@@
Another line of work researches the combination of high-fidelity physics-based models with data-driven models. We identify four distinct approaches in the context of our research: assistant, residual, surrogate, and augmentation. In the assistant strategy, the output from a physics-based model is used as an additional input to the data-driven model [6, 2] or conversely, the data-driven model can provide inputs or corrections to the physics-based model [42, 4, 21]. This additional input may provide valuable context information, but also increases the number of input features. The residual strategy involves using a data-driven model to learn the residuals between the output of a physics-based model and actual observed data [37, 10, 13, 25]. This approach aims to capture unmodeled physical phenomena and variations in the data that the physics-based model may not fully account for. It is particularly useful when certain inputs are not reliably represented in the data or when incorporating domain knowledge with a physical reference input is necessary. In the surrogate strategy, a data-driven model is trained to replace the physics-based model by using the same inputs, with the physics-based model’s simulations serving as the target outputs [45, 44, 15, 41, 35]. The primary motivation for this approach is to reduce computation time, allowing the data-driven model to perform simulations that would otherwise be time-consuming with the physics-based model. Nevertheless, it relies on a sufficiently accurate simulator. The augmentation strategy augments the real data with simulated data from a physics-based model, training the data-driven model on this augmented dataset and subsequently adapting it to real-world situations [27, 9, 1, 40]. This augmentation strategy is particularly advantageous when little to no real-world data is available, as it allows the model to leverage simulated data to improve its accuracy and generalization capabilities. However, this approach may suffer from performance degradation if there is a significant disparity between simulated and real data.
@@
Although hybrid approaches in BEM are gaining increasing attention, three significant research gaps remain. First, most recent studies focus on developing new hybrid methods within specific data contexts, yet there is a lack of a comprehensive research comparing the advantages and disadvantages of different hybrid approaches. A systematic evaluation of these methods is still missing, which limits our understanding of their relative strengths and weaknesses. Second, there are only few works that compare scenarios with limited building documentation or sensor data. Third, the majority of existing literature primarily evaluates hybrid models in terms of accuracy, with less attention given to the explainability of these approaches. Greater emphasis on explainability is crucial for understanding the hybrid models general behaviour, uncovering model biases of the physics-based sub-model and building trust for real-world application.
These research gaps highlight the need for more comparative studies and a deeper exploration of how hybrid BEMs perform under varying levels of data availability and quantity. In our study, we focus on the thermal modeling aspect of the BEM to predict indoor temperature dynamics, as understanding these dynamics forms the foundation for further analysis. Furthermore, we concentrate on a high-fidelity physics-based model for the construction of hybrid models. This paper tries to address the previously mentioned research gaps by making the following three main contributions:
@@ 2 Methodology
Our methodology consists of combining a physics-based EnergyPlus model with a data-driven model in four prevalent combinations. We then proceed to investigate these combinations in terms of accuracy with standard accuracy metric, interpretability with hierarchical Shapley values and data dependency with different documentation and sensor availability scenarios.
@@ 2.1 Hybrid model
In our research, we evaluate the four most prevalent forms of combining physics-based and data-driven models in building energy modeling, shown in Figure 1. The assistant approach incorporates the indoor temperature predictions from the physics-based model as an additional input to the data-driven model. The residual approach involves using a data-driven model to learn the indoor temperature residuals – the differences between the physics-based model’s predictions and the actual indoor temperatures. The surrogate approach aims to train a data-driven model to fully replace the physics-based model for indoor temperature prediction, taking the physics-based simulation as training label. Finally, the augmentation approach takes this step further by fine-tuning the data-driven model with real data following its initial pre-training as a surrogate.
Figure 1: Overview of the different hybrid approaches with physics-based EnergyPlus and data-driven combination. The data sources are indicated as building documentation and sensors. The corresponding data flows are indicated by colored arrows, which meanings are given in the legend at the figure bottom. The orange exogenous data arrow indicates all the sensor measurements apart from the room temperature. $X$ and $Y$ denote the input and target variables associated with the input data and output data of the data-driven model for learning purposes. Note that for the augmentation approach, the $Y^{\prime}$ is used as first learning on simulated data and $Y$ as the fine-tuning step.
@@ 2.2 Physics-based models
As our physics-based model, we utilize the widely recognized open-access building simulation software EnergyPlus (EP) [39] to construct a comprehensive building energy model, as depicted in Figure 2. EnergyPlus simulates the building and its subsystems using a system of differential equations. These modules include equations for the building envelope, indoor heat balance, mass balance, HVAC fluid dynamics, lighting system, window performance, and weather conditions. To set-up the EP model, the building geometry and physical parameters are curated from the building documentation and provided in an Intermediate Data Format file. Alongside weather data formatted in the EnergyPlus weather file, these inputs allow the internal simulation manager to accurately simulate the building’s dynamics behavior. Once the multi-zone building model is set up, its parameters are calibrated by aligning the real target variable with the simulation output using the sensor data from the building and individual rooms. This procedure ensures accurate simulation of temperature dynamics within the building. In case of less detailed building information, an archetypal EnergyPlus model can be constructed. For this purpose, we use the archetype framework Cesar-P [43] for automated model construction.
@@ 2.3 Data-driven models
In the case of data-driven models we investigate Linear Regression (LR), Feedfoward Neural Network (FFNN) and Random Forest (RF). We choose the LR due to its simplicity as well as strong baseline performance. The FFNN is chosen for its popularity and adequacy for regression problems. It also has computational advantages compared to kernel-based methods such as Support Vector Regression. The RF is selected as a representative of the tree ensemble models that are widely used in regression problems for their accuracy. Moreover, both FFNN and RF are able to represent the building temperature dynamics of all rooms within a single model. A Recurrent Network architecture is not considered since no time-lagged features are used as input. This enables a direct comparison to the pure physics-based EnergyPlus simulation, which doesn’t allow explicit time lag integration.
@@
$\mathbf{W}=\operatorname*{arg\,min}_{\mathbf{W}}\dfrac{1}{n}\sum_{i+1}^{N}||\mathbf{y}_{i}-\mathbf{f}(\mathbf{x}_{i},\mathbf{W})||^{2}$ (7)
@@ 2.4 Explainability
In a combination of physics-based and data-driven models, the explainability of the data-driven model as well as of the interdependencies between the two sub-models is crucial for the overall understanding. In particular, we want to examine these two aspects through the lens of feature contributions for model predictions on a global level. In this way, we can assign feature importance to get a better understanding of the model’s general behavior, uncover model biases and potentially give suggestions whether to record particular sensor variables. To comprehensively explain the hybrid models, we employ a Shapley value framework [34] with the SHAP package [23]. Given the challenge of correlated input features, we opt for a hierarchical Shapley value calculation [28]. In this approach, hierarchical Shapley values are recursively computed based on a predefined hierarchy. We establish this hierarchy using agglomerative clustering [33], with Pearson’s distance [16] as the distance metric. Consequently, the distance matrix $\mathbf{D}$ can be expressed as follows:
@@
where the norm ||$\cdot$|| ⋅ | in this context denotes the size of a set, $M$ is the set of clusters, $B=\{B_{1},...,B_{m}\}$ represents the partition of features into clusters, $B_{k}$ is the cluster containing feature $i$ and $k$ is the index of cluster containing feature $i$. $R$ is a subset of clusters and the first binomial coefficient calculates the corresponding number of ways that to choose $|R|$ subsets out of $|M|-1$ clusters. $T$ is a subset of features and the second binomial coefficient calculates the corresponding number of ways that to choose $|T|$ features within cluster $B_{k}$ from $|B_{k}|-1$ features. $Q=\bigcup_{r\in R}B_{r}$ is a union of features in the clusters defined by $R$, and $v(\cdot)=\mathbb{E}[\mathbf{f}(\mathbf{X})|\mathbf{X}_{(\cdot)}]$ is the expected prediction of a subset. Therefore, the term $v(Q\cup T\cup\{i\})-v(Q\cup T)$ describes the marginal contribution of feature $i$ when added to the union $Q\bigcup T$.

### v2

@@ Abstract
Building energy modeling plays a vital role in optimizing the operation of building energy systems by providing accurate predictions of the building’s real-world conditions. In this context, various techniques have been explored, ranging from traditional physics-based models to data-driven models. Recently, researchers are combining physics-based and data-driven models into hybrid approaches. This includes using the physics-based model output as additional data-driven input, learning the residual between physics-based model and real data, learning a surrogate of the physics-based model, or fine-tuning a surrogate model with real data. However, a comprehensive comparison of the inherent advantages of these hybrid approaches is still missing. The primary objective of this work is to evaluate four predominant hybrid approaches in building energy modeling through a real-world case study, with focus on indoor thermodynamics. To achieve this, we devise three scenarios reflecting common levels of building documentation and sensor availability, assess their performance, and analyze their explainability using hierarchical Shapley values. The real-world study reveals three notable findings. First, greater building documentation and sensor availability lead to higher prediction accuracy for hybrid approaches. Second, the performance of hybrid approaches depends on the type of building room, but the residual approach using a Feedforward Neural Network as data-driven sub-model performs best on average across all rooms. This hybrid approach also demonstrates a superior ability to leverage the simulation from the physics-based sub-model. Third, hierarchical Shapley values prove to be an effective tool for explaining and improving hybrid models while accounting for input correlations.
@@ 1 Introduction
In the last decade, the operation of buildings accounted for approximately 30% of global energy consumption and 26% of CO2 emissions [1]. Over half of this consumption stems from the operation of building energy systems such as Heating, Ventilation and Air Conditioning (HVAC), as well as the electrical systems. This situation underscores the urgent need for measures to reduce energy consumption in the operation these systems. Fortunately, advancements in digitalization and sensor deployment provide a foundation for using analytical and machine learning tools to optimize the performance of building energy systems. On this basis, Building Energy Models (BEM) can be developed to simulate a building’s real-world condition and predict future behaviour thereby enabling the recommendation of optimal control actions. However, a major challenge remains: the availability of building information and sensor data, particularly in terms of comprehensive building documentation, sensor coverage, and the length of recorded data. This lack of building information and sensor data complicates the reliable and accurate construction of BEMs but also hinders the widespread adoption of energy optimization strategies in buildings.
The Building Energy Model is a digital representation of a building designed for simulating energy and thermodynamics. It is constructed based on various characteristics of the building, including building geometry, material properties, installed energy systems, and operational inputs such as weather conditions, HVAC operation, and occupancy schedules [2]. To accurately reflect real-world conditions, the BEM is initially calibrated with sensor data to ensure faithful representation of the building and its energy systems. Once calibrated, the BEM can predict building operational performance and indoor temperature evolution across different time scales such as quarter-hourly or hourly resolution. Its comprehensive predictive capabilities make the BEM a versatile tool for various applications in building planning and operations. In retrofitting studies, it serves as an evaluation tool to assess potential measures for improving energy efficiency [3]. In building operations, it provides detailed temperature predictions as basis for control algorithms, enabling the optimization of energy system operations [4].
@@
More recent approaches of building energy modeling focus on data-driven methods, which utilize statistical and machine learning techniques such as Autoregressive Integrated Moving Average (ARIMA), Feedforward Neural Network (FFNN), Long Short-Term Memory (LSTM) and Convolutional Neural Network (CNN) [2]. These methods rely on sensor measurements to establish relationships between defined input and output variables. Commonly modelled relationships include those between building operation data – such as energy consumption or temperature – and the corresponding sensor readings. Since these models are applied to directly data, their underlying mechanisms are not easily accessible, classifying them as data-driven models. Furthermore, their complexity reduces interpretability. In such cases, post-hoc explainability methods like Shapley values [6], can be used to analyze model predictions. Unlike physics-based models, data-driven approaches do not require detailed physical building information or the calibration of physical parameters, making them easier to set up as digital twins. However, their performance is highly dependent on the quantity and quality of the available measurement data. Despite this dependency, data-driven models have demonstrated high accuracy in numerous studies and are particularly well-suited for modeling buildings at an urban scale, thanks to their reduced configuration time [7].
@@
Several research studies aim to combine reduced-order physics-based models with data-driven models through two general strategies. The first strategy involves formulating a loss function based on the reduced-order model to train the data-driven model. A prominent example is the use of Physics-informed Neural Network (PINN), where a physics-informed loss function is devised based on RC model thermodynamics in order to train a FFNN [8, 9, 10]. The second strategy consists of incorporating reduced-order model elements into a data-driven model architecture. A number of works model building dynamics with a State-Space model framework and parameterize state and/or observation equation with FFNN [11], Neural Ordinary Differential Equations [12] or Graph Neural Networks [13]. However, ensuring physical consistency still remains a challenge in aforementioned strategies [14].
@@
Another line of work researches the combination of high-fidelity physics-based models with data-driven models. We identify four distinct approaches in the context of our research: assistant, residual, surrogate, and augmentation. In the assistant strategy, the output from a physics-based model is used as an additional input to the data-driven model [15, 16] or conversely, the data-driven model can provide inputs or corrections to the physics-based model [17, 18, 19]. This additional input may provide valuable context information, but also increases the number of input features. The residual strategy involves using a data-driven model to learn the residuals between the output of a physics-based model and actual observed data [20, 21, 22, 23]. This approach aims to capture unmodelled physical phenomena and variations in the data that the physics-based model may not fully account for. It is particularly useful when certain inputs are not reliably represented in the data or when incorporating domain knowledge with a physical reference input is necessary. In the surrogate strategy, a data-driven model is trained to replace the physics-based model by using the same inputs, with the physics-based model’s simulations serving as the target outputs [24, 25, 26, 27, 28]. The primary motivation for this approach is to reduce computation time, allowing the data-driven model to perform simulations that would otherwise be time-consuming with the physics-based model. Nevertheless, it relies on a sufficiently accurate simulator. The augmentation strategy augments the real data with simulated data from a physics-based model, training the data-driven model on this augmented dataset and subsequently adapting it to real-world situations [29, 30, 31, 32]. This augmentation strategy is particularly advantageous when little to no real-world data is available, as it allows the model to leverage simulated data to improve its accuracy and generalization capabilities. However, this approach may suffer from performance degradation if there is a significant disparity between simulated and real data. An overview of the different characteristics of hybrid approaches in the context of our research is shown in Figure 1.
@@
Although hybrid approaches in BEM are gaining increasing attention, three significant research gaps remain. First, most recent studies focus on developing new hybrid methods within specific data contexts, yet there is a lack of a comprehensive research comparing the advantages and disadvantages of different hybrid approaches. A systematic evaluation of these methods is still missing, which limits our understanding of their relative strengths and weaknesses. Second, there are only few works that compare scenarios with limited building documentation or sensor data. Third, the majority of existing literature primarily evaluates hybrid models in terms of accuracy, with less attention given to the explainability of these approaches. Greater emphasis on explainability is crucial for understanding the hybrid models general behaviour, uncovering model biases of the physics-based sub-model and building trust for real-world application. Moreover, due to environmental and operational physical dependencies, sensor measurements exhibits interaction effects that are reflected in the model features. As a consequence, the joint influence of feature groups and non-linear interactions between features impact model predictions.
Figure 1: Overview of the different hybrid approaches with physics-based EnergyPlus and data-driven combination. The data sources are indicated as building documentation and sensors with various sensor groups. The corresponding data flows are indicated by colored arrows, which meanings are given in the legend at the figure bottom. The green exogenous data arrow indicates all the sensor measurements apart from the indoor room temperature. $X$ and $Y$ denote the input features and target variables of the data-driven model for learning purposes. $\hat{Y}$ represents the hybrid model’s indoor temperature prediction. Note that the augmentation approach involves a two-step learning procedure. While step 1 indicates the learning on simulated data, step 2 denotes the fine-tuning on real indoor temperature.
These research gaps highlight the need for more comparative studies and a deeper exploration of how hybrid BEMs perform under varying levels of data availability and quantity. We address these research gaps by conducting a comprehensive study of hybrid approaches in a real-world setting. For explainability, we opt for post-hoc explanation of model predictions, given the complexity of the investigated models. We introduce the hierarchical Shapley value as a suitable framework for our setting, as it effectively assigns feature contributions to correlated feature groups within the original output unit. In our study, we focus on the thermal modeling aspect of the BEM to predict indoor thermodynamics for two reasons. Firstly, the understanding of thermodynamics forms the foundation for all energy related calculations, also in physics-based models. Secondly, the focus on thermodynamics allows for a clear effect modeling and facilitated model analysis. Furthermore, we concentrate on a high-fidelity physics-based model for the construction of hybrid models. This paper tries to address the previously mentioned research gaps by making the following three main contributions:
@@ 2 Methodology
Our methodology consists of combining a physics-based EnergyPlus model with a data-driven model in four prevalent combinations. We then proceed to evaluate these combinations based on accuracy using standard accuracy metrics, explainability using hierarchical Shapley values, and data dependency under different documentation and sensor availability scenarios.
@@ 2.1 Hybrid model
In our research, we evaluate the four most prevalent forms of combining physics-based and data-driven models in building energy modeling, shown in Figure 1. The assistant approach incorporates the indoor temperature predictions from the physics-based model as an additional input to the data-driven model. The residual approach involves using a data-driven model to learn the indoor temperature residuals – the differences between the physics-based model’s predictions and the real indoor temperatures. The surrogate approach aims to train a data-driven model to fully replace the physics-based model for indoor temperature prediction, taking the physics-based simulation as training label. Finally, the augmentation approach takes this step further by fine-tuning the data-driven model with real data following its initial pre-training as a surrogate.
The corresponding algebraic equations for these hybrid models are given as:
$\mathbf{f}_{assistant}(\mathbf{x})$ $=\mathbf{f}_{dd}(\mathbf{x},\mathbf{f}_{EP}(\mathbf{x})),\quad l_{dd}(\mathbf{y},\mathbf{f}_{assistant}(\mathbf{x}))$ (1)
$\mathbf{f}_{residual}(\mathbf{x})$ $=\mathbf{f}_{dd}(\mathbf{x},\mathbf{f}_{EP}(\mathbf{x})),\quad l_{dd}(\mathbf{y}-\mathbf{f}_{EP}(\mathbf{x}),\mathbf{f}_{residual}(\mathbf{x}))$ (2)
$\mathbf{f}_{surrogate}(\mathbf{x})$ $=\mathbf{f}_{dd}(\mathbf{x},\mathbf{f}_{EP}(\mathbf{x})),\quad l_{dd}(\mathbf{f}_{EP}(\mathbf{x}),\mathbf{f}_{surrogate}(\mathbf{x}))$ (3)
$\mathbf{f}_{augmentation}(\mathbf{x})$ $=\mathbf{f}_{surrogate}(\mathbf{x},\mathbf{f}_{EP}(\mathbf{x})),\quad l_{dd}(\mathbf{y},\mathbf{f}_{augmentation}(\mathbf{x}))$ (4)
where $\mathbf{f}_{EP}$ and $\mathbf{f}_{dd}$ are the physics-based EnergyPlus and data-driven prediction functions, $\mathbf{x}$ is a sample with the corresponding feature dimension, $\mathbf{y}$ is the target with real indoor temperatures, and $l_{dd}$ is the data-driven loss function. We employ all four approaches in our case study and compare them extensively.
@@ 2.2 Physics-based models
As our physics-based model, we utilize the widely recognized open-access building simulation software EnergyPlus (EP) [5] to construct a comprehensive building energy model, as depicted in Figure 2. EnergyPlus simulates the building and its subsystems using a system of differential equations. These modules include equations for the building envelope, indoor heat balance, mass balance, HVAC fluid dynamics, lighting system, window performance, and weather conditions. To set-up the EP model, the building geometry and physical parameters are curated from the building documentation and provided in an Intermediate Data Format file. Alongside weather data formatted in the EnergyPlus weather file, these inputs allow the internal simulation manager to accurately simulate the building’s dynamics behaviour. Once the multi-zone building model is set up, its parameters are calibrated by aligning the real target variable with the simulation output using the sensor data from the building and individual rooms. This procedure ensures accurate simulation of thermodynamics within the building. In case of less detailed building information, an archetypal EnergyPlus model can be constructed. For this purpose, we use the archetype framework Cesar-P [39, 40] for automated model construction. On the basis of building footprint, height, year of construction and building type, the Cesar-P framework constructs a suitable archetypal Energyplus model with pre-defined material and geometry settings.
@@ 2.3 Data-driven models
In the case of data-driven models we investigate Linear Regression (LR), Feedfoward Neural Network (FFNN) and Random Forest (RF). We choose the LR due to its simplicity as well as strong baseline performance. The FFNN is chosen for its popularity and adequacy for regression problems. It also has computational advantages compared to kernel-based methods such as Support Vector Regression. The RF is selected as a representative of the tree ensemble models that are widely used in regression problems for their accuracy. Moreover, both FFNN and RF are able to represent the building thermodynamics of all rooms within a single model. A Recurrent Network architecture is not considered since no time-lagged features are used as input. This enables a direct comparison to the pure physics-based EnergyPlus simulation, which doesn’t allow explicit time lag integration.
@@
$\mathbf{W}=\operatorname*{arg\,min}_{\mathbf{W}}\dfrac{1}{N}\sum_{i=1}^{N}||\mathbf{y}_{i}-\mathbf{f}(\mathbf{x}_{i},\mathbf{W})||^{2}$ (11)
@@ 2.4 Explainability
We refer to explainability as the ability to describe and justify a model’s predictions, especially for complex data-driven models, such as Deep Learning models used in this research [42]. In a combination of physics-based and data-driven models, the explainability of the data-driven model as well as of the interdependencies between the two sub-models is crucial for the overall understanding. In particular, we want to examine these two aspects through the lens of feature contributions for model predictions on a global level. In this way, we can assign feature importance to get a better understanding of the model’s general behaviour, uncover model biases and potentially give suggestions whether to record particular sensor variables. To comprehensively explain the hybrid models, we employ a Shapley value framework [43] with the Shapley additive explanations (SHAP) package [6]. Given the challenge of correlated input features, we opt for a hierarchical Shapley value calculation [44]. In this approach, hierarchical Shapley values are recursively computed based on a predefined hierarchy. We establish this hierarchy using agglomerative clustering [45], with Pearson’s distance [46] as the distance metric. Consequently, the distance matrix $\mathbf{D}$ can be expressed as follows:
@@
where the norm ||$\cdot$|| ⋅ | in this context denotes the size of a set, $M$ is the set of clusters, $B=\{B_{1},...,B_{m}\}$ represents the partition of features into clusters, $B_{k}$ is the cluster containing feature $i$ and $k$ is the index of cluster containing feature $i$. $R$ is a subset of clusters and the first binomial coefficient calculates the corresponding number of ways that to choose $|R|$ subsets out of $|M|-1$ clusters. $T$ is a subset of features and the second binomial coefficient calculates the corresponding number of ways that to choose $|T|$ features within cluster $B_{k}$ from $|B_{k}|-1$ features. $Q=\bigcup_{r\in R}B_{r}$ is a union of features in the clusters defined by $R$, and $v(\cdot)=\mathbb{E}[\mathbf{f}(\mathbf{X})|\mathbf{X}_{(\cdot)}]$ is the expected prediction of a subset. Therefore, the term $v(Q\cup T\cup\{i\})-v(Q\cup T)$ describes the marginal contribution of feature $i$ when added to the union $Q\bigcup T$. In addition to the hierarchical Shapley values, we analyze the intrinsic interpretability of different models: Linear Regression through regression coefficients, Random Forest through Gini importance, and Neural Network through Jacobian matrix sensitivity. However, these intrinsic measures become less interpretable in the presence of feature interactions from correlated input features.